OpenStudy (anonymous):

Factor polynomial completely x(y-9)- 5(y-9)

OpenStudy (anonymous):

(x-5)(y-9)

OpenStudy (anonymous):

(x-9)(y-5)

hero (hero):

ab + ac = a(b+c) x(y-9)- 5(y-9) = (y-9)(x-5)

OpenStudy (anonymous):

I thought it was xy- 9x -5y+45 for some reason